OpenStudy (anonymous):

vanessa built a rectangular pen for her dogs. She used an outside wall of the garage for one of the sides of the pen. She had to buy 20 \(m\) of fencing in order to build the other sides of the pen if its area is 48 \(m^2\) i will include a illustration in the comments.

OpenStudy (anonymous):



OpenStudy (superdavesuper):

Let one side be x as labeled in the drawing n the other length be y, we know 2x + y = 20 What else do we know?

OpenStudy (anonymous):

\(x\times y=48\)

OpenStudy (superdavesuper):

yup so we have 2x + y = 20 and x * y = 48 can we solve for x and y now?

OpenStudy (anonymous):

humm...

OpenStudy (anonymous):

OHHHHH i got it now thanks

OpenStudy (superdavesuper):

hahahaa u are SMART! :)

OpenStudy (anonymous):

:P yea i was just confused :P

OpenStudy (anonymous):

thanks!